# Theoretical Benefit and Limitation of Diffusion Language Model

### v1

@@ Appendix A Auxiliary Lemma
Lemma A.1 (Upper Bound for Multi-tokens Sampling).
@@
Lemma A.2 (Chernoff Bound).
@@
Lemma A.3 (Pinsker’s Inequality).
@@
Appendix B Proof for Theorem 4.2
@@
  • We establish a connection between the perplexity of the discrete diffusion model and the KL divergence between generated and data distributions. This involves deriving an upper bound on KL divergence using expected divergence over reverse processes (Lemma B.2) and decomposing this divergence into per-step conditional KL terms (Lemma B.3).
  • We analyze n-gram model dependencies through a rigorous characterization of reverse processes (Definition B.1) and separators—$(n-1)$ continuous sampled tokens that create independent intervals (Definition B.4). This leads to a precise formulation of per-step dependencies using these separators (Definition B.5).
  • We derive an upper bound for the KL divergence between generated and data distributions based on the number of per-step dependencies (Lemmas B.9 and B.8).
  • We employ probabilistic bounds to analyze and bound the number of per-step dependencies (Lemmas B.10, B.11 and A.1).
  • Finally, we demonstrate the existence of a schedule achieving small KL divergence with $O(\frac{n-1}{\epsilon^{n}})$ steps by constructing an efficient sampling schedule using the preceding lemmas (Lemma B.12).
@@
Definition B.1 (An Instance of Reverse Process).
@@
Lemma B.2 (KL Divergence Upper Bound for the Masking Schedule).
@@
Lemma B.3 (KL Divergence Decomposition for the Reverse Process).
@@
Definition B.4 (Number of Separators in a Reverse Step).
@@
Definition B.5 (Number of Dependencies in a Reverse Step).
@@
Example B.6 (Computing Dependencies in the $n$-gram Setting).
@@
Definition B.7 (Number of Dependencies in a Reverse Process).
@@
Lemma B.8 (KL Divergence Upper Bound for the Instance of Reverse Process).
@@
Using Lemma B.3, we have:
@@
By Lemma A.1, we know that:
@@
Lemma B.9 (KL Divergence Upper Bound for a Masking Schedule).
@@
Applying Lemma B.8 to the instances of reverse process, we can conclude that:
@@
Lemma B.10 (Bounds on Separator and New Location Count at Each Reverse Step).
@@
Applying Lemma A.2 to the sum of $X_{j}$, we derive:
@@
Since the sampling of each location occurs independently, applying Lemma A.2, we have:
@@
Lemma B.11 (Upper Bound for the Expectation of Dependencies at Each Reverse Step).
@@
By Lemma B.10, at step $t_{i}$, the following bounds hold:
@@
Using the same trick as Lemma B.10, applying Lemma A.2, we can derive the bound for probability $\Pr\left(\operatorname{NEW}\geq x\right)$ as:
@@
Lemma B.12 (Efficient Sampling with Small KL Divergence).
@@
By Lemma B.9, we know that:
@@
Note that at step $t_{1}$, the reverse process can be bounded using Lemma A.1. By reviewing our proof process, it is easy to see that we can substitute $\operatorname{DEP}_{n}({\mathcal{M}}_{1},{\mathcal{M}}_{<1})$ for $(|{\mathcal{M}}_{1}|-1)\log|{\mathcal{V}}|$, where ${\mathcal{V}}$ stands for the vocabulary. By the definition of $\delta_{i}$, we know that:
@@
Applying Lemma B.11 to $\operatorname{DEP}_{n}({\mathcal{M}}_{i},{\mathcal{M}}_{<i})$, if $L\delta_{i}\geq 1$, we can obtain:
@@
Theorem B.13 ($\operatorname{TER}$ Bounds for $n$-Gram Language Generation).
@@
By Lemma B.12, for any $L>O\big{(}\frac{n-1}{\epsilon^{n+0.5}}\big{)}$, there exists a masking schedule $\alpha_{t}$ with $N=O\big{(}\frac{n-1}{\epsilon^{n}}\big{)}$ sampling steps satisfying:
@@
Appendix C Proof for Theorem 4.3 and Theorem 4.4
C.1 Proof for Theorem 4.3
@@
Lemma C.1 (Low Probability of Simultaneous Sampling with Sufficient Steps).
@@
By Lemma B.10, we know that the probability of a location being sampled at time step $t_{i}=\frac{N-i}{N}$ is:
@@
Lemma C.2 (Accurate Step-by-Step Generation with Low Learning Error).
@@
Theorem C.3 (Accurate Generation of HMM with Sufficient Steps).
@@
By Lemma C.1, given the masking schedule $\alpha_{t}$, there exists $N_{0}$, for $N>N_{0}$ and $N$ reverse steps, the probability of sampling multiple locations in the sequence at the same time is bounded by:
@@
In other words, the probability of sampling all the locations at different steps is at least $\frac{1-\delta}{e^{-\delta}}$. By Lemma C.2, for each reverse process which satisfies that all the locations are sampled at different steps, the probability of generating a valid sequence is lower bounded by:
@@
C.2 Proof for Theorem 4.4
In the section, we aim to find an example (Example C.7) with high sequence error rate. To present this example, we begin with a special class of languages defined under the interval setting:
Definition C.4 (Interval Setting).
Consider a sequence of length $L$, which is divided equally into $M$ intervals ${\mathcal{I}}_{1},{\mathcal{I}}_{2},\cdots,{\mathcal{I}}_{M}$, each of length $l=\frac{L}{M}\geq 2$. Given a masking schedule $\alpha_{t}$, an instance of reverse process $\tau=({\mathcal{M}}_{1},{\mathcal{M}}_{2},\cdots,{\mathcal{M}}_{N})$ is defined by Definition B.1. For any two locations within different intervals, their corresponding tokens are independent from each other. In other words, let $\tilde{{\bm{x}}}_{i}^{(j)}$ denote the new tokens in ${\mathcal{M}}_{i}\cap{\mathcal{I}}_{j}$, $\tilde{{\bm{x}}}_{<i}^{(j)}$ denote the previously sampled tokens in ${\mathcal{M}}_{<i}\cap{\mathcal{I}}_{j}$, and $p$ denote the distribution over sequences generated by the reverse model with reverse process $\tau$, then for time step $t_{i}=\frac{N-i}{N}$:
@@
Lemma C.5 (Simultaneous Sampling Probability for an Interval).
@@
Let $\delta_{i}=\alpha_{t_{i}}-\alpha_{t_{i-1}}$. Similar to Lemma B.10, we know that $\delta_{i}$ is the probability of a location being sampled at time step $t_{i}$. Take the first location in $|{\mathcal{I}}_{j}|$, denote it as $X_{1}$, and let $X_{2},\cdots,X_{l}$ denote the rest $l-1$ locations in ${\mathcal{I}}_{j}$. If $X_{1}$ is sampled at step $t_{i}$, then $X_{2},\cdots,X_{l}$ must be sampled at time steps other than $t_{i}$. Therefore, $h_{j}$ can be bounded by:
@@
Lemma C.6 ($\operatorname{SER}$ bound for Interval Setting).
@@
By Lemma C.5, we can obtain that for each interval ${\mathcal{I}}_{j}$, the probability $p_{\textrm{error}}^{(j)}$ of generating an error in ${\mathcal{I}}_{j}$ is lower-bounded by:
@@
In this setup, each interval ${\mathcal{I}}_{k}$ contains $\frac{l(l-1)}{2}$ pairs of mutually independent random variables. Given an arbitrary masking schedule $\alpha_{t}$, this setting is consistent with Definition C.4. Let $q$ denote the data distribution described above.
Under 4.1, we only need to examine the case where ${\bm{x}}_{t}$ has no error. By Lemma A.3, we know that:
@@
In other words, this is consistent with the setting Lemma C.6, with an error probability $p_{e}=\frac{1}{2}-2\sqrt{2{\epsilon}_{\textit{learning}}}$.
@@
Note C.8 (HMM Form of Example C.7).
The setting described in Example C.7 can be alternatively modeled as a Hidden Markov Model (HMM), where the observation space is ${\mathcal{O}}={\mathcal{A}}$, and the state space is ${\mathcal{S}}=\{(i,A^{(i)})|A^{(i)}\in\mathbb{R}^{(l-1)\times(l-1)},i=1,\cdots,l\}$. Here, $i$ represents the current position within the interval, and $A^{(i)}$ is an upper triangular matrix with entries taking values of 0 or 1. For $j\leq i$, the $j$-th row of $A^{(i)}$ encodes the values sampled by the variable pairs formed between the $j$-th position and all its subsequent positions in the interval. For $j>i$, the $j$-th row of $A^{(i)}$ is set to 0.
@@
The above Note gives the HMM form of Example C.7. In fact, with appropriate adjustments, it can be further modified into an n-gram language. Using the HMM defined above, we can prove Theorem 4.4.
Theorem C.9 (SER Bound for HMM Generation).
@@
Take the HMM described in Note C.8, and set $l=5$, $N=CL$. The vocabulary is the observation space ${\mathcal{O}}$ which satisfies $|{\mathcal{O}}|=2^{l-1}$. By Lemma C.6, for any masking schedule $\alpha_{t}$, we have:
@@
Appendix D Extend to Efficient Sampling Strategies
  1. 1.
  2. 2.
Appendix E Experiment Details

### v2

@@ Appendix C Auxiliary Lemma
Lemma C.1 (Upper Bound for Multi-tokens Sampling).
@@
Lemma C.2 (Chernoff Bound).
@@
Lemma C.3 (Pinsker’s Inequality).
@@
Appendix D Proof for Theorem 4.2
@@
  • We establish a connection between the perplexity of the discrete diffusion model and the KL divergence between generated and data distributions. This involves deriving an upper bound on KL divergence using expected divergence over reverse processes (Lemma D.2) and decomposing this divergence into per-step conditional KL terms (Lemma D.3).
  • We analyze n-gram model dependencies through a rigorous characterization of reverse processes (Definition D.1) and separators—$(n-1)$ continuous sampled tokens that create independent intervals (Definition D.4). This leads to a precise formulation of per-step dependencies using these separators (Definition D.5).
  • We derive an upper bound for the KL divergence between generated and data distributions based on the number of per-step dependencies (Lemmas D.9 and D.8).
  • We employ probabilistic bounds to analyze and bound the number of per-step dependencies (Lemmas D.10, D.11 and C.1).
  • Finally, we demonstrate the existence of a schedule achieving small KL divergence with $O(\frac{n-1}{\epsilon^{n}})$ steps by constructing an efficient sampling schedule using the preceding lemmas (Lemma D.12).
@@
Definition D.1 (An Instance of Reverse Process).
@@
Lemma D.2 (KL Divergence Upper Bound for the Masking Schedule).
@@
Lemma D.3 (KL Divergence Decomposition for the Reverse Process).
@@
Definition D.4 (Number of Separators in a Reverse Step).
@@
Definition D.5 (Number of Dependencies in a Reverse Step).
@@
Example D.6 (Computing Dependencies in the $n$-gram Setting).
@@
Definition D.7 (Number of Dependencies in a Reverse Process).
@@
Lemma D.8 (KL Divergence Upper Bound for the Instance of Reverse Process).
@@
Using Lemma D.3, we have:
@@
By Lemma C.1, we know that:
@@
Lemma D.9 (KL Divergence Upper Bound for a Masking Schedule).
@@
Applying Lemma D.8 to the instances of reverse process, we can conclude that:
@@
Lemma D.10 (Bounds on Separator and New Location Count at Each Reverse Step).
@@
Applying Lemma C.2 to the sum of $X_{j}$, we derive:
@@
Since the sampling of each location occurs independently, applying Lemma C.2, we have:
@@
Lemma D.11 (Upper Bound for the Expectation of Dependencies at Each Reverse Step).
@@
By Lemma D.10, at step $t_{i}$, the following bounds hold:
@@
Using the same trick as Lemma D.10, applying Lemma C.2, we can derive the bound for probability $\Pr\left(\operatorname{NEW}\geq x\right)$ as:
@@
Lemma D.12 (Efficient Sampling with Small KL Divergence).
@@
By Lemma D.9, we know that:
@@
Note that at step $t_{1}$, the reverse process can be bounded using Lemma C.1. By reviewing our proof process, it is easy to see that we can substitute $\operatorname{DEP}_{n}({\mathcal{M}}_{1},{\mathcal{M}}_{<1})$ for $(|{\mathcal{M}}_{1}|-1)\log|{\mathcal{V}}|$, where ${\mathcal{V}}$ stands for the vocabulary. By the definition of $\delta_{i}$, we know that:
@@
Applying Lemma D.11 to $\operatorname{DEP}_{n}({\mathcal{M}}_{i},{\mathcal{M}}_{<i})$, if $L\delta_{i}\geq 1$, we can obtain:
@@
Theorem D.13 ($\operatorname{TER}$ Bounds for $n$-Gram Language Generation).
@@
By Lemma D.12, for any $L>O\big{(}\frac{n-1}{\epsilon^{n+0.5}}\big{)}$, there exists a masking schedule $\alpha_{t}$ with $N=O\big{(}\frac{n-1}{\epsilon^{n}}\big{)}$ sampling steps satisfying:
@@
Appendix E Proof for Theorem 4.3 and Theorem 4.4
E.1 Proof for Theorem 4.3
@@
Lemma E.1 (Low Probability of Simultaneous Sampling with Sufficient Steps).
@@
By Lemma D.10, we know that the probability of a location being sampled at time step $t_{i}=\frac{N-i}{N}$ is:
@@
Lemma E.2 (Accurate Step-by-Step Generation with Low Learning Error).
@@
Theorem E.3 (Accurate Generation of HMM with Sufficient Steps).
@@
By Lemma E.1, given the masking schedule $\alpha_{t}$, there exists $N_{0}$, for $N>N_{0}$ and $N$ reverse steps, the probability of sampling multiple locations in the sequence at the same time is bounded by:
@@
In other words, the probability of sampling all the locations at different steps is at least $\frac{1-\delta}{e^{-\delta}}$. By Lemma E.2, for each reverse process which satisfies that all the locations are sampled at different steps, the probability of generating a valid sequence is lower bounded by:
@@
E.2 Proof for Theorem 4.4
In the section, we aim to find an example (Example E.7) with high sequence error rate. To present this example, we begin with a special class of languages defined under the interval setting:
Definition E.4 (Interval Setting).
Consider a sequence of length $L$, which is divided equally into $M$ intervals ${\mathcal{I}}_{1},{\mathcal{I}}_{2},\cdots,{\mathcal{I}}_{M}$, each of length $l=\frac{L}{M}\geq 2$. Given a masking schedule $\alpha_{t}$, an instance of reverse process $\tau=({\mathcal{M}}_{1},{\mathcal{M}}_{2},\cdots,{\mathcal{M}}_{N})$ is defined by Definition D.1. For any two locations within different intervals, their corresponding tokens are independent from each other. In other words, let $\tilde{{\bm{x}}}_{i}^{(j)}$ denote the new tokens in ${\mathcal{M}}_{i}\cap{\mathcal{I}}_{j}$, $\tilde{{\bm{x}}}_{<i}^{(j)}$ denote the previously sampled tokens in ${\mathcal{M}}_{<i}\cap{\mathcal{I}}_{j}$, and $p$ denote the distribution over sequences generated by the reverse model with reverse process $\tau$, then for time step $t_{i}=\frac{N-i}{N}$:
@@
Lemma E.5 (Simultaneous Sampling Probability for an Interval).
@@
Let $\delta_{i}=\alpha_{t_{i}}-\alpha_{t_{i-1}}$. Similar to Lemma D.10, we know that $\delta_{i}$ is the probability of a location being sampled at time step $t_{i}$. Take the first location in $|{\mathcal{I}}_{j}|$, denote it as $X_{1}$, and let $X_{2},\cdots,X_{l}$ denote the rest $l-1$ locations in ${\mathcal{I}}_{j}$. If $X_{1}$ is sampled at step $t_{i}$, then $X_{2},\cdots,X_{l}$ must be sampled at time steps other than $t_{i}$. Therefore, $h_{j}$ can be bounded by:
@@
Lemma E.6 ($\operatorname{SER}$ bound for Interval Setting).
@@
By Lemma E.5, we can obtain that for each interval ${\mathcal{I}}_{j}$, the probability $p_{\textrm{error}}^{(j)}$ of generating an error in ${\mathcal{I}}_{j}$ is lower-bounded by:
@@
In this setup, each interval ${\mathcal{I}}_{k}$ contains $\frac{l(l-1)}{2}$ pairs of mutually independent random variables. Given an arbitrary masking schedule $\alpha_{t}$, this setting is consistent with Definition E.4. Let $q$ denote the data distribution described above.
Under 4.1, we only need to examine the case where ${\bm{x}}_{t}$ has no error. By Lemma C.3, we know that:
@@
In other words, this is consistent with the setting Lemma E.6, with an error probability $p_{e}=\frac{1}{2}-2\sqrt{2{\epsilon}_{\textit{learning}}}$.
@@
Note E.8 (HMM Form of Example E.7).
The setting described in Example E.7 can be alternatively modeled as a Hidden Markov Model (HMM), where the observation space is ${\mathcal{O}}={\mathcal{A}}$, and the state space is ${\mathcal{S}}=\{(i,A^{(i)})|A^{(i)}\in\mathbb{R}^{(l-1)\times(l-1)},i=1,\cdots,l\}$. Here, $i$ represents the current position within the interval, and $A^{(i)}$ is an upper triangular matrix with entries taking values of 0 or 1. For $j\leq i$, the $j$-th row of $A^{(i)}$ encodes the values sampled by the variable pairs formed between the $j$-th position and all its subsequent positions in the interval. For $j>i$, the $j$-th row of $A^{(i)}$ is set to 0.
@@
The above Note gives the HMM form of Example E.7. In fact, with appropriate adjustments, it can be further modified into an n-gram language. Using the HMM defined above, we can prove Theorem 4.4.
Theorem E.9 (SER Bound for HMM Generation).
@@
Take the HMM described in Note E.8, and set $l=5$, $N=CL$. The vocabulary is the observation space ${\mathcal{O}}$ which satisfies $|{\mathcal{O}}|=2^{l-1}$. By Lemma E.6, for any masking schedule $\alpha_{t}$, we have:
@@
E.3 Extending Theorem 4.4 to Remasking Strategies
In this section, we extend the conclusion of Theorem 4.4 to masked diffusion models with remasking sampler (named ReMDM) proposed by (Wang et al., 2025). We begin by presenting the main mechanisms of the sampler, followed by the justification for analogous results.
As is introduced in the paper, ReMDM is an MDM with remasking designs, where preciously sampled tokens have a chance to be remasked and resampled again in later steps. Formally, let $\sigma_{t}$ be the remasking schedule that satisfies $0\leq\sigma_{t}\leq\min\{1,\frac{1-\alpha_{s}}{\alpha_{t}}\}$ for any time step $s<t$.
In contrast to the conventional reverse process defined in Equation 2, the reverse process of ReMDM (with the original sequence ${\bm{x}}$) is defined as:
$\begin{gathered}q_{s|t}({\bm{x}}_{s}|{\bm{x}}_{t},{\bm{x}})=\prod_{i=0}^{L-1}q_{s|t}(x_{s}^{i}|{\bm{x}}_{t},{\bm{x}}),\quad\text{where}\\ q_{s|t}(x_{s}^{i}|{\bm{x}}_{t},{\bm{x}})=\begin{cases}1-\sigma_{t},&x_{t}^{i}\neq\texttt{[m]},x_{s}^{i}=x_{t}^{i},\\ \sigma_{t},&x_{t}^{i}\neq\texttt{[m]},x_{s}^{i}=\texttt{[m]},\\ \frac{1-\alpha_{s}-\sigma_{t}\alpha_{t}}{1-\alpha_{t}},&x_{t}^{i}=\texttt{[m]},x_{s}^{i}=\texttt{[m]},\\ \frac{\alpha_{s}-(1-\sigma_{t})\alpha_{t}}{1-\alpha_{t}}q_{0|t}(x_{s}^{i}|{\bm{x}}_{t}),&x_{t}^{i}=\texttt{[m]},x_{s}^{i}\neq\texttt{[m]},\\ 0,&\text{otherwise.}\end{cases}\end{gathered}$ (7)
Intuitively, compared to the original reverse process, there is a probability that already generated tokens are remasked during the generation. Now, consider the last time $t$ that a location is sampled. In other words, the location is sampled at step $t$, and is not remasked in later steps, thus the token sampled at $t$ will stay unchanged. We will derive the probability that a location is sampled for the last time at time step $t_{i}=\frac{N-i}{N}$.
Lemma E.10 (Last Sample Probability Estimate of ReMDM).
Given a masking schedule $\alpha_{t}$, a remasking schedule $\sigma_{t}$ and $N$ reverse steps, for ReMDM, the probability $\delta_{i}$ that a location is sampled for the last time at time step $t_{i}=\frac{N-i}{N}$ satisfies that:
$\delta_{i}=(\alpha_{t_{i}}-(1-\sigma_{t_{i-1}})\alpha_{t_{t-1}})\prod_{j=i}^{N-1}(1-\sigma_{t_{j}}).$
Proof.
It follows from the symmetry that we can consider any location in the sequence. First, we use induction to show the probability $p_{m}(i)$ that the token at this location is [m]at $t_{i}$ satisfies $p_{m}(i)=1-\alpha_{t_{i}}$.
For $i=0$, since $\alpha_{1}=0$, it is direct that $p_{m}(0)=1=1-\alpha_{t_{0}}$.
For $i=k+1$, assume that $p_{m}(k)=1-\alpha_{t_{k}}$. Combining with Equation 7, we know that
$p_{m}(k+1)=p_{m}(k)\cdot\frac{1-\alpha_{t_{k+1}}-\sigma_{t_{k}}\alpha_{t_{k}}}{1-\alpha_{t_{k}}}+(1-p_{m}(k))\cdot\sigma_{t_{k}}=1-\alpha_{t_{k+1}}.$
Thus, we can conclude that $p_{m}(i)=1-\alpha_{t_{i}}$. Therefore, the probability $\delta_{i}$ can be decomposed into
$\delta_{i}$ $=\Pr(\text{{[m]}at $t_{t_{i-1}}$})\cdot\Pr(\text{not {[m]}at $t_{t_{i}}$}\mid\text{{[m]}at $t_{t_{i-1}}$})\cdot\Pr(\text{stay unmasked}\mid\text{not {[m]}at $t_{t_{i}}$})$
$=p_{m}(i-1)\cdot\frac{\alpha_{t_{i}}-(1-\sigma_{t_{i-1}})\alpha_{t_{i-1}}}{1-\alpha_{t_{i-1}}}\cdot\prod_{j=i}^{N-1}(1-\sigma_{t_{j}})$
$=(\alpha_{t_{i}}-(1-\sigma_{t_{i-1}})\alpha_{t_{t-1}})\prod_{j=i}^{N-1}(1-\sigma_{t_{j}}).$
By replacing the probability of a location being sampled at time step $t$ with the probability of that a location being sampled for the last time to be at time step $t$ (and letting $\delta_{i}$ to be the corresponding probability in Lemma E.10), we can derive analogous results to those presented in Lemma E.5, Lemma E.6, and Theorem E.9. This substitution is justified because when two locations are sampled simultaneously and are not later remasked and resampled, there exists a probability of introducing errors. Therefore, by applying similar proof techniques, we arrive at the same conclusion as stated in Theorem 4.4.
Appendix F Experiment Details